# Detecting local perturbations of networks in a latent hyperbolic embedding space

### v2

@@ Abstract
Keywords Brain Networks  $\cdot$ Geometric embedding  $\cdot$ Epilepsy
@@ 1 Introduction
Any system whose behavior can be seen as resulting from the interaction of a set of interconnected units can be modeled in the mathematical form of a graph. The last two decades has seen a growing interest in applying graph theory to describe the complex structure and function of real-world interconnected systems, such as the Internet, social networks, and the brain. Evidence has revealed [1] that these systems share some peculiar characteristics, like the small-world property [2], hierarchical organization [3], and efficient navigability [4]. Therefore, understanding and analyzing them would require an expansion of standard graph models towards more specific and advanced models and methods.
@@
Network geometry, particularly hyperbolic geometry, has recently gained attention as a plausible latent geometric model of complex networks, from which typical properties such as a heterogeneous degree distribution and strong clustering can naturally be induced by the metric structure [16]. At the same time, the approach of projecting the network onto a geometric space, called network embedding, has been used to produce highly informative low-dimensional representations of networks. Embedding networks in Euclidean spaces is a current approach to studying and comparing their connectivity structure [17]. However, such embeddings often require the use of high dimensions and fail to encompass some graph topological properties, such as its hierarchical structure. In contrast, non-Euclidean mapping methods often require a lower number of dimensions to accurately embed graphs and better reflect their organization [18].
@@ 2.1 Network embedding in hyperbolic spaces
Hyperbolic embedding is a non-linear method that takes into account the hierarchical structure of the graph (if any), and depicts both the hierarchies and the similarity of the nodes in the embedding [22]. Embedding in hyperbolic space replicates properties of the graph, such as sparseness, self-similarity, scale-free degree distribution, small-world characteristics, and community structure, that embedding in Euclidean spaces cannot capture [23]. Hyperbolic embeddings have been proven to be useful for community detection and link prediction, as the embedding captures the similarity between network nodes in a graph [24, 23]. Therefore, we chose hyperbolic space, instead of Euclidean space, for the network embedding. This decision allowed us to capture the properties of the original network in a lower-dimensional space and quantitatively determine local perturbations to the connectivity structure.
@@ 2.2 Perturbation detection from network embeddings
$\cosh\mathrm{dist}_{hyp}(i,j)=\cosh r_{i}\times\cosh r_{j}\\ -\sinh r_{i}\times\sinh r_{j}\times\cos(\pi-|\pi-|\theta_{i}-\theta_{j}||)$ (1)
@@
We define HypDisp as the displacement of a node in the hyperbolic disk caused by a perturbation. To estimate this, we embed both the initial and perturbed networks in hyperbolic space using the technique described above in Section § 2.1. Next, we align the two embeddings, as very often, a small perturbation may introduce a random angular offset to the nodes’ positions in the hyperbolic disk. A weak perturbation may thus result in two embeddings with the same similarities between nodes but with a different structure regarding the coordinates [28]. The alignment is achieved by rotating the perturbed embedding until the minimum sum of HypDisp, calculated between all the nodes of the whole perturbed and original embedded networks, is achieved. After aligning both embeddings, a hyperbolic displacement is calculated for each node (see Figure 1). For each node $i$, its positions resulting from the original and perturbed networks are $Pos_{o}(i)$ and $Pos_{p}(i)$, respectively. The score $\mathrm{HypDisp}(i)$ attributed to the node $i$ is then :
@@
To compute the hyperbolic distortion, we consider the same group of nodes in the embedding relative to the perturbed network $G^{\prime}$, denoted as $kN^{\prime}(i)$, and quantify the change in their hyperbolic distances from $i^{\prime}$ due to the perturbation. We call $i^{\prime}$ the node $i$ in the perturbed network $G^{\prime}$, and we use it to specify that the set of nodes nearest to $i^{\prime}$ in the embedding space of $G^{\prime}$, $kN^{\prime}(i^{\prime})$, is generally different from $kN^{\prime}(i)$. We note that $kN(i)$ and $kN^{\prime}(i)$ indicate the same network nodes, but in two different embedding configurations, meaning they will generally have different coordinates. These concepts are illustrated on the right side of Figure 1.
Let’s explain how the score works in the specific: for each node $i$ we compute the set of distances $\{\mathrm{dist}_{hyp}(i,j),j\in kN(i)\}$ in the embedding space of the reference network $G$. We compare these values to $\{\mathrm{dist}_{hyp}(i^{\prime},j),j\in kN^{\prime}(i)\}$ by estimating the score function:
$\mathrm{kNHypD}(i)={{\sum}}_{j\in kN^{\prime}(i)}\frac{\left|\mathrm{dist}_{hyp}(i^{\prime},j)-\mathrm{dist}_{hyp}(i,j)\right|}{\mathrm{dist}_{hyp}(i,j)}$ (3)
@@
The best choice of $k$ in the estimation of kNHypD depends upon the data. As for other non-parametric estimators in statistics [29], the variance of the estimator appears to be high for small values of $k$. Although increasing the size of the neighborhood reduces the variance, it may also yield very coarse estimates of kNHypD thus increasing the bias. Moreover, the stability gained by using higher values of $k$ can be overshadowed by the complexities involved in computing a large number of distances. Overall, based on the estimator’s performance across the different networks considered, it is recommended to choose a neighborhood size between 10
@@ 3.2 Clinical data
Finally, we assessed our scores using anatomical brain connectivity networks derived from patients who underwent epilepsy surgery. We examined the brain networks of 51 patients, estimated using T1-weighted structural and diffusion-weighted MRI scans before and after surgery. The pre- and post-operative networks were obtained from an open dataset associated with the research paper by Sinha et al. [12]. The dataset included 51 individuals who underwent epilepsy surgery, from which 30 patients underwent surgery for a left temporal lobe epilepsy (TLE), and 21 for a right temporal lobe epilepsy. Within each category, participants were further divided based on post-surgical results: those with positive surgical outcomes (seizure-free after surgery) and those with negative outcomes (not seizure-free after surgery). The surgical procedure involved removing brain tissue located in the presumed epileptic zone (specifically, in the temporal lobe) to prevent seizures. In graph theory, this corresponds to cutting a subset of edges within a localized section of the network. The brain networks before the surgery have the following properties: a mean degree of $28\pm 2$ nodes, a standard deviation of the degree distribution of $12\pm 1$, a mean clustering coefficient of $0.57\pm 0.02$. After surgery, the perturbed networks display a mean degree of $25\pm 2$, a standard deviation of the degree distribution of $11\pm 1$, and a clustering coefficient of $0.55\pm 0.03$. Brain networks’ size is $N=114$ nodes and there are no disconnected nodes after surgery.
@@ 4.1 Synthetic networks
For both network models, the scores calculated from the second-order neighbors of targeted nodes are very low and similar to those obtained from the baseline, suggesting a good localization power, as these nodes are practically unaffected by perturbations. Strong local correlations of nodes (assortativity), however, could have an effect and induce a bias in the scores of these “second-order target nodes”. Interestingly, at the average degree kNHypD scores show a peak in Watts-Strogatz networks and a crossing with the baseline in the Barabási-Albert networks. In contrast, no clear trend was found for the kNHypD score as a function of the target node’s clustering coefficient (results in the Supplementary Material). Therefore, we can conclude that: i) the proposed perturbation scores have good localization and a small bias; ii) the performance improves with perturbation magnitude; and iii) the performances of the kNHypD score depend on degree heterogeneity.
@@ 4.2 Brain networks
To illustrate the effectiveness of our scores on real data and compare their performances, we used them on brain network connectivity perturbations due to epilepsy surgery (refer to §3.2 for further details). For illustration purposes, here we considered patients with a positive surgical outcome. For these patients, we compared two networks: the pre- and post-surgery networks. The network after surgery is a modified version of the original network, with fewer edges specifically removed in the surgery area. A way to assess the perturbation of node $i$ due to surgery is the perturbation degree, $P_{deg}(i)$, defined as $P_{deg}(i)=\nicefrac{{(D_{pre}(i)-D_{post}(i))}}{{D_{pre}(i)}}$, where $D_{pre}(i)$ and $D_{post}(i)$ denote the pre- and post-surgery degree of the ithnode. When $P_{deg}(i)\rightarrow 0$, the node’s connectivity is slightly affected by the surgery. In contrast, when $P_{deg}(i)\rightarrow 1$, the node is more likely to become disconnected and significantly impacted.
@@
Figure 6: Left and top views of brain network perturbations in epileptic patients who have undergone left hemisphere surgery. Each node’s size corresponds to the surgery perturbation rate (rate of removed connections) of the node. The color of nodes represents its score by either HypDisp (score 1) or kNHypD (score 2).
@@ Acknowledgements
Authors are grateful to Steven Rico for useful and valuable discussions during the preparation of the study. A.L. was supported by the doctoral school EDITE from Sorbonne University. M.G. acknowledges doctoral support from the Ecole Normale Supérieure. We thank the two anonymous referees for helpful comments.
@@
The brain network data that support the findings of this study are available are freely and publicly available in the supplementary material of this article [12].

### v3

@@ 1 Introduction
Any system whose behavior can be seen as resulting from the interaction of a set of interconnected units can be modeled in the mathematical form of a graph. The last two decades have seen a growing interest in applying graph theory to describe the complex structure and function of real-world interconnected systems, such as the Internet, social networks, and the brain. Evidence has revealed [1] that these systems share some peculiar characteristics, like the small-world property [2], hierarchical organization [3], and efficient navigability [4]. Therefore, understanding and analyzing them would require an expansion of standard graph models towards more specific and advanced models and methods.
@@
Network geometry, particularly hyperbolic geometry, has recently gained attention as a plausible latent geometric model of complex networks, from which typical properties such as a heterogeneous degree distribution and strong clustering can naturally be induced by the metric structure [16]. At the same time, the approach of projecting the network onto a geometric space, called network embedding, has been used to produce highly informative low-dimensional representations of networks. Embedding networks in Euclidean spaces is a current approach to study and compare their connectivity structure [17]. However, such embeddings often require the use of high dimensions and fail to encompass some graph topological properties, such as its hierarchical structure. In contrast, non-Euclidean mapping methods often require a lower number of dimensions to accurately embed graphs and better reflect their organization [18].
@@ 2.1 Network embedding in hyperbolic spaces
Hyperbolic embedding is a non-linear method that takes into account the hierarchical structure of the graph (if any), and depicts both the hierarchies and the similarity of the nodes in the embedding [22]. Embedding in hyperbolic space replicates properties of the graph, such as sparseness, self-similarity, scale-free degree distribution, small-world characteristics, and community structure, that embedding in Euclidean spaces cannot capture [23]. Hyperbolic embeddings have been proven to be useful for community detection and link prediction, as the embedding captures the similarity between network nodes in a graph [24, 23]. Therefore, we chose hyperbolic space, instead of Euclidean one, for the network embedding. This decision allowed us to capture the properties of the original network in a lower-dimensional space and quantitatively determine local perturbations to the connectivity structure.
@@ 2.2 Perturbation detection from network embeddings
$\cosh\mathrm{dist}_{hyp}(i,j)=\cosh r_{i}\times\cosh r_{j}-\sinh r_{i}\times\sinh r_{j}\times\cos(\pi-|\pi-|\theta_{i}-\theta_{j}||)$ (1)
@@
We define HypDisp as the displacement of a node in the hyperbolic disk caused by a perturbation. To estimate this, we embed both the initial and perturbed networks in hyperbolic space using the technique described above in Section § 2.1. Next, we align the two embeddings, as very often, a small perturbation may introduce a random angular offset to the nodes’ positions in the hyperbolic disk. A weak perturbation may thus result in two embeddings with the same similarities between nodes but with a different structure regarding the coordinates [28]. The alignment is achieved by rotating the perturbed embedding until the minimum sum of HypDisp, calculated between all the nodes of the whole perturbed and original embedded networks, is reached. After aligning both embeddings, a hyperbolic displacement is calculated for each node (see Figure 1). For each node $i$, its positions resulting from the original and perturbed networks are $Pos_{o}(i)$ and $Pos_{p}(i)$, respectively. The score $\mathrm{HypDisp}(i)$ attributed to the node $i$ is then :
@@
To compute the hyperbolic distortion, we consider the same group of nodes in the embedding relative to the perturbed network $G^{\prime}$, denoted as $kN(i^{\prime})$, and quantify the change in their hyperbolic distances from $i^{\prime}$ due to the perturbation. We call $i^{\prime}$ the node corresponding to $i$ in the perturbed network $G^{\prime}$. We note that $kN(i)$ and $kN(i^{\prime})$ indicate the same set of nodes, but in two different embedding configurations, meaning they will generally have different coordinates. We also specify that the set of nodes nearest to $i^{\prime}$ in the embedding space of $G^{\prime}$, $kN^{\prime}(i^{\prime})$, is generally different from $kN(i^{\prime})$. These concepts are illustrated on the right side of Figure 1.
Let’s explain how the score works in the specific: for each node $i$ we compute the set of distances $\{\mathrm{dist}_{hyp}(i,j),j\in kN(i)\}$ in the embedding space of the reference network $G$. We compare these values to $\{\mathrm{dist}_{hyp}(i^{\prime},j),j\in kN(i^{\prime})\}$ by estimating the score function:
$\mathrm{kNHypD}(i)={{\sum}}_{j\in kN(i)}\frac{\left|\mathrm{dist}_{hyp}(i^{\prime},j)-\mathrm{dist}_{hyp}(i,j)\right|}{\mathrm{dist}_{hyp}(i,j)}$ (3)
@@
The best choice of $k$ in the estimation of kNHypD depends upon the data. As for other non-parametric estimators in statistics [29], the variance of the estimator appears to be high for small values of $k$. Although increasing the size of the neighborhood reduces the variance, it may also yield very coarse estimates of kNHypD thus increasing the bias. Moreover, the stability gained by using higher values of $k$ can be overshadowed by the complexities involved in computing a large number of distances. Overall, based on the score’s performances across the different networks considered, it is recommended to choose a neighborhood size between 10% and 50% of the network size.
@@ 3.2 Clinical data
Finally, we assessed our scores using anatomical brain connectivity networks derived from patients who underwent epilepsy surgery. We examined the brain networks of 51 patients, estimated using T1-weighted structural and diffusion-weighted MRI scans before and after surgery. The pre- and post-operative networks were obtained from an open dataset associated with the research paper by Sinha et al. [12]. The dataset included 51 individuals who underwent epilepsy surgery, from which 30 patients underwent surgery for a left temporal lobe epilepsy (TLE), and 21 for a right temporal lobe epilepsy. Within each category, participants were further divided based on post-surgical results: those with positive surgical outcomes (seizure-free after surgery) and those with negative outcomes (not seizure-free after surgery). The surgical procedure involved removing brain tissue located in the presumed epileptic zone (specifically, in the temporal lobe) to prevent seizures. In graph theory, this corresponds to cutting a subset of edges within a localized section of the network. The brain networks before the surgery have the following properties: a mean degree of $28\pm 2$ nodes, a standard deviation of the degree distribution of $12\pm 1$, a mean clustering coefficient of $0.57\pm 0.02$. After surgery, the perturbed networks display a mean degree of $25\pm 2$, a standard deviation of the degree distribution of $11\pm 1$, and a clustering coefficient of $0.55\pm 0.03$. Brain networks have $N=114$ nodes and there are no disconnected nodes after surgery.
@@ 4.1 Synthetic networks
For both network models, the scores calculated from the second-order neighbors of targeted nodes are similar to those obtained from the baseline, suggesting a good localization power, as these nodes are practically unaffected by perturbations. Strong local correlations of nodes (assortativity), however, could have an effect and induce a bias in the scores of these “second-order target nodes”. Interestingly, at the average degree kNHypD scores show a peak in Watts-Strogatz networks and a crossing with the baseline in the Barabási-Albert networks. In contrast, no clear trend was found for the kNHypD score as a function of the target node’s clustering coefficient (results in the Supplementary Material). Therefore, we can conclude that: i) the proposed perturbation scores have good localization and a small bias; ii) the performance improves with perturbation magnitude; and iii) the performances of the kNHypD score depend on degree heterogeneity.
@@ 4.2 Brain networks
To illustrate the effectiveness of our scores on real data and compare their performances, we used them on brain network connectivity perturbations due to epilepsy surgery (refer to §3.2 for further details). For illustration purposes, here we considered patients with a positive surgical outcome. For these patients, we compared two networks: the pre- and post-surgery networks. The network after surgery is a modified version of the original network, with fewer edges specifically removed in the surgery area. A way to assess the perturbation of node $i$ due to surgery is the perturbation degree, $P_{deg}(i)$, defined as $P_{deg}(i)=\nicefrac{{(D_{pre}(i)-D_{post}(i))}}{{D_{pre}(i)}}$, where $D_{pre}(i)$ and $D_{post}(i)$ denote the pre- and post-surgery degree of the ithnode. When $P_{deg}(i)\rightarrow 0$, the node’s connectivity is slightly affected by the surgery. In contrast, when $P_{deg}(i)\rightarrow 1$, the node is more likely to become disconnected and significantly impacted. In both groups (left and right hemisphere surgery), three nodes out of the 114 have a perturbation magnitude greater than $0.8$, while 109 nodes have a perturbation degree below $0.4$. The average perturbation magnitude across all patients and nodes is $0.118\pm 0.032$.
@@
Figure 6: Left and top views of brain network perturbations in epileptic patients who have undergone left hemisphere surgery. Each node’s size corresponds to the surgery perturbation rate (rate of removed connections) of the node. The color of nodes represents its score by either HypDisp (score 1) or kNHypD (score 2). For the sake of visualisation, values are normalized between 00 and $1$.
@@ Acknowledgements
Authors are grateful to Steven Rico for useful and valuable discussions during the preparation of the study. A.L. was supported by the doctoral school EDITE from Sorbonne University. M.G. acknowledges doctoral support from the Ecole Normale Supérieure de Lyon. We thank the two anonymous referees for helpful comments.
@@
The brain network data that support the findings of this study are publicly available in the supplementary material of this article [12].